# PAD: Phase-Amplitude Decoupling Fusion for Multi-Modal Land Cover Classification

### v1

@@ III-B Overall Framework
As illustrated in Fig. 2, the proposed PAD framework integrates SAR and RGB modalities via a dual-branch encoder and a shared decoder, with its primary innovation in cross-modal feature fusion through phase-amplitude decoupling. The encoder utilizes asymmetric convolutional backbones $\mathcal{B}_{\mathrm{SAR}}$ and $\mathcal{B}_{\mathrm{RGB}}$ with independent parameters to preserve modality-specific attributes. Specifically, $\mathcal{B}_{\mathrm{SAR}}$ extracts microwave scattering features from SAR images, while $\mathcal{B}_{\mathrm{RGB}}$ captures optical reflectance patterns from RGB images. Both backbones produce hierarchical feature maps at four stages with resolutions $\{1/4,1/8,1/16,1/32\}$, such that $i\in\{0,1,2,3\}$:
@@
Initially, the SCF module (Eq. 5) processes the bimodal features by computing pixel-wise fusion weights through spatial attention gating. This operation preliminarily integrates complementary features in the spatial domain and reduces feature dimensionality, thereby constraining the solution space for subsequent frequency-domain operations.
@@
The fused features are then transformed via FD into separate amplitude and phase spectra:
@@
The amplitude spectrum $\mathcal{A}^{i}$ is enhanced by the ASF module, which distinguishes high-frequency details from low-frequency structures using a trainable frequency radius $r$:
$\mathcal{A^{\prime}}^{i}=\mathrm{ASF}\left(\mathcal{A}^{i};r\right).$ (7)
Simultaneously, the phase spectrum is corrected via multiplicative scaling in the PSC module:
@@
Finally, FR reconstructs the features, yielding the final fusion output:
@@
The shared decoder, $\mathcal{D}$, aggregates the multi-stage fused features to generate high-precision land-cover segmentation maps:
@@ III-C Frequency Decoupling & Frequency Recoupling
To achieve physically disentangled representations in the spectral domain, we propose a bidirectional transformation framework comprising FD and its inverse, FR. First, the input image $X\in\mathbb{R}^{H\times W}$ is transformed from the spatial domain to the frequency domain using a 2D real-valued fast Fourier transform (rFFT) [49], yielding the spectrum $\mathcal{F}\in\mathbb{C}^{H\times W}$:
$\mathcal{F}\left(u,v\right)=\sum_{x=0}^{W-1}\sum_{y=0}^{H-1}X\left(x,y\right)\exp\left(-i2\pi\left(\frac{ux}{W}+\frac{vy}{H}\right)\right),$ (11)
where $(u,v)$ represents frequency-domain coordinates. Exploiting the conjugate symmetry of real-valued signals, only the non-redundant region (with $20\leq u\leq W/$ 2 and $0\leq v<H$) is retained. Next, an FFT Shift operation relocates low-frequency components to the center of the spectrum:
@@
where $\oplus$ denotes modulo addition. Finally, the amplitude and phase spectra are derived as:
@@
with $\mathcal{A}\in\mathbb{R}^{H\times W}$ capturing radiometric properties and $\mathcal{P}\in\mathbb{R}^{H\times W}$ preserving structural details. This decoupling enables separate learning of SAR scattering features and RGB structural priors.
@@ III-D Phase Spectrum Correction Module
As illustrated in Fig. 1(c), minor misalignments in the low-frequency phase spectra between SAR and RGB images, attributable to inherent geometric distortions (e.g., side-looking SAR vs. RGB orthographic projection), can compromise spatial consistency and boundary precision. Traditional phase alignment methods relying on rigid transformations are inadequate for non-rigid distortions. Thus, we propose a lightweight PSC module that performs adaptive geometric correction using learnable multiplicative masks. Given an input phase spectrum tensor $P\in\mathbb{R}^{C\times H\times W}$, the module first compresses the channel dimension via a $1\times 1$ convolution (reducing $C$ to $C/r$, where $r\in\{4,8\}$) and applies a Leaky ReLU activation [50]. A subsequent $1\times 1$ convolution produces a single-channel Phase Correction Mask (PCM) constrained to $(0,1)$ by a Sigmoid function. Pixel-wise multiplicative scaling is then applied as follows:
$P^{\prime}=P\odot\left(1+\text{PCM}\right),$ (14)
where $\odot$ denotes the Hadamard product. This design limits the maximum scaling factor to 2, ensuring the preservation of phase periodicity while requiring significantly fewer parameters than conventional $3\times 3$ convolutions.
@@ III-E Amplitude Spectrum Fusion Module
Our analysis in Fig. 1(b) indicates that the spectral discrepancy between SAR and RGB modalities is primarily concentrated in the medium-to-high frequency components of the amplitude spectra. To exploit this prior while maintaining gradient stability, we design an adaptive spectrum fusion module with three sequential operations. First, we compute a normalized frequency distance metric:
@@
which quantifies component positions relative to the spectrum center $(X_{\mathrm{center}},Y_{\mathrm{center}})$. A trainable boundary radius $r$, constrained via a Sigmoid activation to yield $R=\sigma(r)$, dynamically partitions the spectrum into low- and high-frequency regions. To maintain differentiability, a soft partition is employed:
@@
with temperature parameter $\tau=10$. The high-frequency components are then enhanced using a position-wise multilayer perceptron (MLP):
@@
and concatenated with the original amplitude spectra. A subsequent $1\times 1$ convolution and residual connection yield the fused amplitude:
@@
This module preserves low-frequency structures while adaptively reinforcing high-frequency details that capture complementary semantic information. The learnable radius $R$ self-adjusts to the input characteristics, eliminating the need for manual threshold tuning.
@@ III-F2 Auxiliary Semantic Supervision
An auxiliary segmentation head is added at the second-to-last stage ($t=2$) using an identical cross-entropy loss:
@@ III-F3 Phase-Amplitude Decoupling Constraint
To resolve the coupling between amplitude and phase spectra, we introduce an Amplitude Consistency Loss that regularizes the fused amplitude $\mathcal{A}^{\prime}_{t}$ to remain faithful to its original decoupled state $\mathcal{A}_{t}$ across feature scales:
@@
where $\mathcal{T}=\{0,1,2,3\}$ indexes the feature scales, $\alpha_{t}=0.5+(1.0-0.5)\times\frac{t}{3}$ imposes depth-progressive weighting, and $\|\cdot\|_{F}$ is the Frobenius norm.
@@ IV-A Dataset Description
WHU-OPT-SAR [23] is a multi-modal dataset developed for land use classification in Hubei Province, China. The dataset comprises 100 pairs of RGB, near-infrared (NIR), and SAR images covering identical geographic regions, with corresponding land type labels. The RGB images are acquired from GF1 satellite data, while the SAR images are sourced from GF3 satellite data. Both image types have been resampled to a uniform spatial resolution of 5 meters, with dimensions of 5556 $\times$ 3704 pixels. The dataset spans seven primary land use categories: farmland, urban areas, villages, water bodies, forests, roads, and others. In our study, images are cropped into non-overlapping patches of 512 $\times$ 512 pixels, resulting in 7000 total patches; 80% (5600 patches) are used for training and 20% (1400 patches) for testing.
DDHR-SK is a subset of the DDHR multi-modal dataset developed for land cover classification (LCC) in Pohang City, South Korea [24]. RGB images are collected from GF2 satellite data and SAR images from GF3 satellite data. To simulate adverse conditions such as clouds and fog, the RGB images were processed with a fog filter using the GNU Image Manipulation Program (GIMP). The dataset comprises 6173 paired cloud-affected RGB and SAR images, resampled to a spatial resolution of 1 meter and resized to 256 $\times$ 256 pixels. It covers five land cover types: buildings, farmland, greenery, water, and roads, with a public split of 3087 training images and 3086 validation images.
@@ IV-B Training Setting
All models were trained and evaluated on a single NVIDIA H800 GPU, while the comparative experiments were carried out on a single NVIDIA A100 GPU. Training was conducted using the MMSegmentation framework [51] with a batch size of 4 and 80k iterations, including validation every 8k iterations. We employed the AdamW [52] optimizer with an initial learning rate of $1\times 10^{-4}$ and a Cosine Annealing schedule (with T_max set to the maximum iteration count). A fixed random seed of 42 was used to ensure reproducibility.
Evaluation metrics include mean intersection over union (mIoU), Kappa coefficient, overall accuracy (OA), and F1 score, with performance reported at the iteration yielding the best mIoU. To assess robustness across diverse scenarios and mitigate class imbalance bias, the Kappa coefficient is computed for each sample, and the mean value is reported:
@@
where $P_{o}^{(i)}$ and $P_{e}^{(i)}$ denote the observed and expected agreements for the $i$-th sample, respectively.
@@ IV-C Experimental Results on Datasets
To assess the effectiveness of the proposed PAD method, we compared its performance with 17 different models, including 9 state-of-the-art multi-modal segmentation approaches and 8 cutting-edge unimodal segmentation methods. All experimental results were obtained through our replication of the original experimental configurations. Experiments were performed on two datasets: WHU-OPT-SAR, the largest multi-modal LCC dataset, and DDHR-SK, which contains RGB images captured under cloudy conditions.
TABLE II: Experiment Results (%) on WHU-OPT-SAR.
@@ IV-C1 Evaluations on WHU-OPT-SAR
As summarized in Table II, our PAD framework achieves significant performance improvements on WHU-OPT-SAR. Compared with the best unimodal method (DANet-RGB [13], achieving 53.25% mIoU and 83.19% OA) and the multi-modal baseline MRFS [25] (55.53% mIoU, 84.21% OA), PAD establishes a new state-of-the-art with 56.26% mIoU, 84.56% OA, 53.93% mKappa, and 70.14% mF1. In fine-grained categories, PAD attains 43.24% IoU on Road and 49.70% on Village segmentation, outperforming MRFS by 1.98% and 0.27%, respectively, thereby underlining the precision of our cross-modal fusion strategy.
Figure 6: Visualization Results on WHU-OPT-SAR. Four representative cases are shown: (a) building-dense area, (b) specular water bodies, (c) vegetated region, and (d) urban–rural transition zone. For each case, the following images are presented from left to right: RGB image, SAR image, pseudo-color ground truth, and prediction maps from six competing methods. PAD is compared against leading single-modal baselines (DANet-RGB, TransUNet-SAR) as well as multi-modal approaches (CDDFuse, CMX, MRFS). Regions of interest are highlighted with gray boxes (for building/water areas) and circles (for vegetation zones); further details are provided in Section IV-C.
Visual analyses support these quantitative results. For example, in Fig. 6(a), PAD successfully recovers occluded road segments (highlighted in cyan) in building-dense regions through SAR phase analysis, while RGB-only models ignore these areas. In Fig. 6(b), PAD effectively distinguishes specular reflection artifacts from actual water boundaries by leveraging SAR’s dielectric properties, overcoming limitations inherent in optical sensors under cloud interference. Additionally, Fig. 6(c) demonstrates PAD’s ability to resolve spectral ambiguities in vegetation classification, and Fig. 6(d) shows that PAD maintains cross-category consistency in complex urban–rural transitions.
TABLE III: Experiment Results (%) on DDHR-SK.
@@
TABLE IV: Ablation Experimental Results (%) on Main Modules on WHU-OPT-SAR.
ID Components IoU per Category aAcc mKappa mF1 mIoU

### v2

@@ III-B Overall Framework
As illustrated in Fig. 2, the proposed PAD framework integrates SAR and RGB modalities via a dual-branch encoder and a shared decoder, with its primary innovation in cross-modal feature fusion through phase-amplitude decoupling. The encoder employs asymmetric convolutional backbones $\mathcal{B}_{\mathrm{SAR}}$ and $\mathcal{B}_{\mathrm{RGB}}$, each with independent parameters to preserve modality-specific attributes. Specifically, $\mathcal{B}_{\mathrm{SAR}}$ extracts microwave scattering features from SAR images, while $\mathcal{B}_{\mathrm{RGB}}$ captures optical reflectance patterns from RGB images. Both backbones produce hierarchical feature maps at four stages with resolutions $\{1/4,1/8,1/16,1/32\}$ for $i\in\{0,1,2,3\}$:
@@
Although frequency-domain fusion is more robust to modality differences, it can discard rich local interactions in the spatial domain. Therefore, we first introduce a lightweight SCF module to retain fine-grained spatial priors. The SCF module (see Eq. 5) processes the bimodal features by computing pixel-wise fusion weights via spatial attention gating. This operation preliminarily integrates complementary features in the spatial domain and reduces feature dimensionality, constraining the solution space for subsequent frequency-domain operations:
@@
The fused feature $X^{i}$ is then transformed via FD into separate amplitude $\mathcal{A}^{i}$ and phase $\mathcal{P}^{i}$ spectra:
@@
The amplitude spectrum $\mathcal{A}^{i}$ is enhanced by the ASF module, which uses a trainable frequency radius $r$ to distinguish and reinforce high-frequency details while preserving low-frequency structures:
$\mathcal{A^{\prime}}^{i}=\mathrm{ASF}\left(\mathcal{A}^{i};r\right),$ (7)
where the PSC module applies multiplicative scaling to correct the phase:
@@
Finally, FR reconstructs the fused features:
@@
The shared decoder $\mathcal{D}$ then aggregates the multi-stage fused features to generate high-precision land-cover segmentation maps:
@@ III-C Frequency Decoupling & Frequency Recoupling
To achieve physically disentangled representations in the spectral domain, we propose a bidirectional transformation framework comprising FD and its inverse, FR. First, the input image $X\in\mathbb{R}^{H\times W}$ is transformed from the spatial domain to the frequency domain using a 2D real-input fast Fourier transform (rFFT) [48], yielding the spectrum $\mathcal{F}\in\mathbb{C}^{H\times(\lfloor W/2\rfloor+1)}$:
$\mathcal{F}\left(u,v\right)=\sum_{y=0}^{H-1}\sum_{x=0}^{W-1}X\left(y,x\right)\exp\left(-i2\pi\left(\frac{ux}{W}+\frac{vy}{H}\right)\right),$ (11)
where $(u,v)$ are frequency-domain coordinates. Exploiting the conjugate symmetry of real-valued signals, we retain only the non-redundant region ($0\leq u\leq\lfloor W/2\rfloor$, $0\leq v<H$). Next, an FFT shift relocates low-frequency components to the spectrum center:
@@
where $\oplus$ denotes modulo addition. Finally, we derive the amplitude and phase spectra as
@@
where $\mathcal{A},\mathcal{P}\in\mathbb{R}^{H\times(\lfloor W/2\rfloor+1)}$ capturing radiometric and structural information, respectively. This decoupling allows the model to focus on modality-specific representations: SAR phase carries scattering patterns sensitive to physical surface characteristics, whereas RGB amplitude captures fine-grained textures and semantic content. By separating them, the network avoids early-stage interference between heterogeneous modalities, enhancing feature discriminability before integration.
@@ III-D Phase Spectrum Correction Module
While previous works have assumed perfect phase alignment [17], our spectral analysis (see Fig. 1(e)) reveals that low-frequency geometric inconsistencies persist, necessitating a learnable, spatially aware correction mechanism. Inspired by channel attention, we propose a lightweight Phase Spectrum Correction (PSC) module tailored for frequency-domain phase alignment.
As illustrated in Fig. 1(e), minor misalignments in the low-frequency phase spectra between SAR and RGB, attributable to geometric distortions (e.g., side-looking SAR vs. RGB orthographic projection), can compromise spatial consistency and boundary precision. Traditional rigid registration methods are inadequate for such non-rigid distortions. Thus, the PSC module performs adaptive geometric correction using learnable multiplicative masks.
Unlike the global channel scaling in traditional attention mechanisms, the proposed PSC estimates a spatially varying modulation mask via a stack of $d$ consecutive $1\times 1$ convolutions, enabling localized, fine-grained phase adjustment. Given an input phase-spectrum tensor $P\in\mathbb{R}^{C\times H\times(\lfloor W/2\rfloor+1)}$, we first compress the channel dimension via $d$ $1\times 1$ convolution (reducing $C$ to $C/r$, $r\in\{4,8\}$) followed by leaky linear rectification function (Leaky ReLU) [49]. A second $1\times 1$ convolution produces a single-channel Phase Correction Mask (PCM), constrained to $(0,1)$ by a Sigmoid:
$\mathrm{PCM}=\sigma\left(\mathrm{Conv}_{d}^{1\times 1}\left(\delta\left(\cdots\delta\left(\mathrm{Conv}_{1}^{1\times 1}\left(P\right)\right)\cdots\right)\right)\right),$ (14)
where $d$ denotes the number of stacked convolutions and $\delta(\cdot)$ is Leaky ReLU. Pixel-wise multiplicative scaling is then applied:
$P^{\prime}=P\odot\left(1+\mathrm{PCM}\right),$ (15)
where $\odot$ is the Hadamard product. This formulation enables localized, fine-grained phase adjustment while preserving phase periodicity and incurs minimal computational overhead.
@@ III-E Amplitude Spectrum Fusion Module
Previous frequency-domain methods often apply fixed high- or low-pass filters or spatially uniform spectral attention, failing to adapt to content-specific frequency distributions. In contrast, our analysis in Fig. 1(d) shows that the spectral discrepancy between SAR and RGB modalities is primarily concentrated in the medium-to-high frequency components of the amplitude spectra. To exploit this prior while maintaining gradient stability, we design the ASF module to employ a trainable frequency radius for dynamic separation and enhancement of complementary frequency bands via three sequential operations.
First, we compute a normalized frequency distance metric for each spectral component $i$:
@@
where quantifies component positions relative to the spectrum center $(X_{\mathrm{center}},Y_{\mathrm{center}})$. A trainable boundary radius $r$, constrained via a Sigmoid activation to yield $R=\sigma(r)\in[0,1]$, dynamically partitions the spectrum into low- and high-frequency regions. This constraint matches the normalized spectral dimensions and prevents $R$ from diverging to unreasonable values during training, which would otherwise lead to unstable partitioning and degraded performance.
To maintain differentiability, we employ a soft partition:
@@
where temperature $\tau=10$. High-frequency components are enhanced via a position-wise MLP:
@@
then concatenated with the original amplitude $A$ and fused by a $1\times 1$ convolution and residual connection:
@@
This module preserves low-frequency structures while adaptively reinforcing high-frequency details that capture complementary semantic information. The learnable radius $R$ self-adjusts to input characteristics, eliminating manual threshold tuning.
@@ III-F2 Auxiliary Semantic Supervision
An auxiliary segmentation head is added at the second-to-last stage ($t=2$), also using cross-entropy loss:
@@ III-F3 Phase-Amplitude Decoupling Constraint
To resolve the coupling between amplitude and phase spectra and enforce consistency between fused and original amplitudes, we introduce an amplitude consistency loss that regularizes $\mathcal{A}^{\prime}_{t}$ to remain faithful to its original decoupled state $\mathcal{A}_{t}$ across feature scales:
@@
where $\mathcal{T}=\{0,1,2,3\}$ indexes the feature scales, $\alpha_{t}=0.5+0.5\frac{t}{3}$ imposes depth-progressive weighting, and $\|\cdot\|_{F}$ is the Frobenius norm.
@@ IV-A Dataset Description
WHU-OPT-SAR [28] is a multi-modal dataset for land use classification in Hubei Province, China. It contains 100 co-registered RGB, NIR, and SAR image triples with corresponding land type labels. RGB images are acquired by GF-1, and SAR images by GF-3; both are resampled to 5 m resolution and sized 5556 $\times$ 3704 pixels. Seven primary land use categories are covered: farmland, urban areas, villages, water bodies, forests, roads, and others. We crop each image into non-overlapping 512 $\times$ 512 patches, yielding 7,000 patches, of which 80% (5,600) are used for training and 20% (1,400) for testing.
DDHR-SK is a subset of the DDHR multi-modal dataset for land cover classification in Pohang City, South Korea [29]. RGB images from GF-2 and SAR images from GF-3 are paired, and RGB images are processed with a fog filter in GIMP to simulate adverse conditions. The dataset comprises 6,173 cloud-affected RGB-SAR pairs, resampled to 1 m resolution and resized to 256 $\times$ 256 pixels. It includes five land cover types: buildings, farmland, greenery, water, and roads, with 3,087 training and 3,086 validation images in the public split.
@@ IV-B Training Setting
All models were trained using the MMSegmentation framework [50] on a single NVIDIA H800 GPU (comparisons used an NVIDIA A100). We used a batch size of 4 for 80 k iterations, validating every 8 k iterations. Optimization was performed with AdamW [51], an initial learning rate of $1\times 10^{-4}$, and a cosine annealing schedule (T_max = total iterations). A fixed seed (42) ensured reproducibility.
We report mean IoU (mIoU), Kappa coefficient, overall accuracy (OA), and F1 score, selecting the iteration with the best mIoU. To mitigate class-imbalance bias, we compute the Kappa coefficient per sample and report its mean:
@@
where $P_{o}^{(i)}$ and $P_{e}^{(i)}$ are the observed and expected agreements for sample $i$.
@@ IV-C Experimental Results on Datasets
To assess the effectiveness of PAD, we replicated the original configurations of 17 comparison models, nine state-of-the-art multi-modal segmentation approaches and eight leading unimodal methods, on both WHU-OPT-SAR and DDHR-SK.
TABLE II: Quantitative Results (%) on WHU-OPT-SAR. We report per-category IoU for seven land cover classes (Farmland, City, Village, Water, Forest, Road, Others), along with OA, mKappa, mF1, and mIoU. Bold and underlined values indicate the best and second-best performance, respectively.
@@ IV-C1 Evaluations on WHU-OPT-SAR
As shown in Table II, PAD achieves new state-of-the-art results: 56.26% mIoU, 84.56% OA, 53.93% mKappa, and 70.14% mF1. This surpasses the best unimodal method (DANet-RGB [18]: 53.25% mIoU, 83.19% OA) and the multi-modal baseline MRFS [30] (55.53% mIoU, 84.21% OA). In fine-grained classes, PAD attains 43.24% IoU on roads and 49.70% on villages, improving over MRFS by 1.98% and 0.27%, respectively.
Figure 6: Visualization Results on WHU-OPT-SAR. Four representative scenarios: (a) building-dense area, (b) specular water bodies, (c) vegetated region, and (d) urban-rural transition. For each scenario, images are shown (left to right): RGB, SAR, pseudo-color ground truth, and prediction maps from six competing methods. We compare PAD against single-modal baselines (DANet-RGB, TransUNet-SAR) and multi-modal approaches (CDDFuse, CMX, MRFS). Regions of interest are highlighted with gray boxes (buildings/water) and circles (vegetation); see Section IV-C for details.
Visual results in Fig. 6 further illustrate PAD’s benefits. In (a), PAD recovers occluded road segments (cyan) in building-dense areas via SAR phase cues, whereas RGB-only models miss them. In (b), PAD distinguishes specular reflections from true water boundaries under cloud interference. In (c)-(d), PAD resolves vegetation spectral ambiguities and maintains consistency across urban-rural transitions.
TABLE III: Quantitative Results (%) on DDHR-SK. We report per-category IoU for seven land cover classes (Buildings, Roads, Greenery, Water, Farmland), along with OA, mKappa, mF1, and mIoU.
@@
IV-C2 Evaluations on DDHR-SK
As shown in Table III, PAD exhibits superior performance under adverse (cloudy) conditions on DDHR-SK. It achieves 94.59% mIoU and 97.65% OA, outperforming leading unimodal methods (e.g., Mask2former-RGB: 90.25% mIoU, 94.76% mF1; DANet-SAR: 83.31% mIoU, 90.28% mF1) and the multi-modal baseline MRFS (93.57% mIoU, 82.24% mKappa). PAD also attains 82.24% mKappa and 96.60% mF1, with boundary-sensitive categories achieving 99.47% IoU on water and 86.12% IoU on roads, surpassing MRFS by 2.44% and 0.08%, respectively. Qualitative results in Fig. 7 further demonstrate PAD’s ability to delineate water infiltration in irrigated fields and maintain continuous road networks beneath fog.
Figure 7: Visualization Results on DDHR-SK. Four boundary-focused scenarios: (a) coastal demarcation, (b) water-cropland interface, (c) vegetation-building complex, and (d) road network topology. For each scenario, panels show the RGB input, SAR coherence map, pseudo-color ground truth, and prediction maps from six methods. We compare PAD against unimodal baselines (Mask2former-RGB, DANet-SAR) and multi-modal approaches (CDDFuse, CMX, MRFS). Critical boundary regions are highlighted with orange boxes; see Section IV-C for details.
The visual results further demonstrate PAD’s capability for cross-modal boundary refinement. In Fig. 7(a), PAD accurately reconstructs periodically submerged coastlines via SAR coherence analysis, whereas RGB-only models confuse tidal sediments with terrestrial areas. In Fig. 7(b), PAD leverages SAR features to delineate water-infiltration boundaries in irrigated fields, overcoming the over-segmentation common in optical sensors. Moreover, in Fig. 7(c) and Fig. 7(d), PAD achieves continuous road networks under vegetation occlusion and precise material boundaries in vegetation-building complexes, respectively.
TABLE IV: Computational Efficiency Comparison. Model size (#Params), inference time per image, and throughput (FPS) for competitive methods. Lower #Params and inference time, and higher FPS, indicate better efficiency.
Method #Params (M) Inference Time (ms) FPS (frames/s)
CMX [31] 15.00 206.70 1,957.8
MRFS [30] 10.30 695.85 1,342.2
ASANet [32] 7.82 72.91 2,245.2
PAD (Ours) 3.88 5.16 3,228.7
IV-C3 Computational Efficiency Analysis
We evaluate the computational efficiency of PAD in terms of model size (Fusion Modules Only), inference time per image, and frames per second (FPS), as summarized in Table IV. We compare against four high-mIoU methods, CMX [31], MRFS [30], ASANet [32], and CDDFuse [12] (excluded from detailed analysis due to its two-stage design and lower efficiency). All metrics are aggregated across fusion stages.
As shown in Table IV, PAD uses only 3.882 M parameters, achieves the fastest inference time (5.16 ms per image), and delivers the highest FPS (3,228.7). These results demonstrate that PAD not only matches or exceeds accuracy of existing methods but also offers real-time efficiency suitable for operational deployment.
IV-C4 Ablation Study
We performed comprehensive ablation studies on WHU-OPT-SAR to validate the contributions of our core components: the Amplitude Spectrum Fusion (ASF) module, the Phase Spectrum Correction (PSC) module, and Multi-Stage Fusion (MSF).
[…]
ID Components IoU per Category OA mKappa mF1 mIoU